1. The perimeter of a rectangle is 120 feet. The width of the rectangle is
equal to one half the length. Find the length and width of the rectangle?
2. The area of a rectangle is 576msqaured. The length of the rectangle is 4 times greater than the width. What are the length and width of the rectangle?

3, One integer is 8 less than another integer. The product of the lesser integer and -5 is 28 more than the greater integer. What are the integers?

Answer by ankor@dixie-net.com(22740)

I'll give you the equations, see if you can do these now.
:
1. The perimeter of a rectangle is 120 feet. The width of the rectangle is
equal to one half the length. Find the length and width of the rectangle?
Let x = width
2x = Length
Perimeter equation
2(2x) + 2x = 120
Find x
:
:
2. The area of a rectangle is 576msqaured. The length of the rectangle is 4 times greater than the width. What are the length and width of the rectangle?
x = the width
4x = the length
Area equation
4x * x = 576
4x^2 = 576
Solve for x
:
3, One integer is 8 less than another integer. The product of the lesser
integer and -5 is 28 more than the greater integer. What are the integers?
x = the larger integer
(x-8) = the lesser integer
Write an equation for the statement:
"The product of the lesser integer and -5 is 28 more than the greater integer."
-5(x-8) = x + 28
Find x
:
:
If you have any questions about these, you can email me, but give it a try.